OpenStudy (anonymous):

PLEASE HELP Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (kirbykirby):

\((x-5)(x+3)(x+1)(x-3i)(x+3i)\) If you know the zero of your polynomial, then you can factor out the polynomial as \(x-x_1)(x-x_2)\cdots\) where \(x_1,x_2, \ldots\) are the zeros. We multiply though the last term \((x-3i)\) by its conjugate though since you need real coefficients, and \((x-3i)(x+3i)\) is a factorization of a difference of squares: \((x^2+9) = (x^2 -(3i)^2)\)

OpenStudy (kirbykirby):

Just expand this product to get it in standard form

OpenStudy (anonymous):

i got x4-4x3+15x2+25+150

OpenStudy (kirbykirby):

It needs to be of degree 5 since you have a product of 5 terms each with x in them.

OpenStudy (anonymous):

does that mean that the exponents add up to five? (degree 5)?

OpenStudy (kirbykirby):

it means that the term with the highest exponent is 5, that means you need some term \(x^5\)

OpenStudy (anonymous):

i dont think thats right... none of the answers i have to choose from are to the degree of 5 then

OpenStudy (anonymous):

f(x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150 f(x) = x4 - 9x2 - 50x - 150

OpenStudy (kirbykirby):

Oooh I see why... I am terribly sorry I misread the zeroes... I read it as 5, -3, -1, 3i ... my bad! ! the product is then \((x-5)(x+3)(x-(-1+3i))(x-(-1-3i))=(x-5)(x+3)(x+1-3i)(x+1+3i) \)

OpenStudy (anonymous):

then how do you do that?

OpenStudy (kirbykirby):

just multiply it out

OpenStudy (anonymous):

i cant see all of what you wrote... :(

OpenStudy (kirbykirby):

\((x-5)(x+3)(x-(-1+3i))(x-(-1-3i))=\\ (x-5)(x+3)(x+1-3i)(x+1+3i)\)

OpenStudy (anonymous):

i dont understand how to multiply the 3i

OpenStudy (kirbykirby):

just treat it as any number that you usually use.